OpenStudy (anonymous):

Subtracting a radical? Please help me!!!!! (7- √54) - (5 + √24)

OpenStudy (anonymous):

\[\sqrt{54}=\sqrt{9*6}\] \[\sqrt{24}=\sqrt{4*6}\]

OpenStudy (anonymous):

use that to simplify... then combine like terms

OpenStudy (anonymous):

2 - √30?

OpenStudy (anonymous):

why sqrt30 ?

OpenStudy (anonymous):

9 is 3^2 so sqrt(9*6) = 3*sqrt(6)

OpenStudy (anonymous):

2 - 5 sqrt 6?

OpenStudy (anonymous):

yes